BOOTSTRAPPING THE CHANGE-POINT OF A HAZARD RATE

TUAN D. PHAM^{1} AND HUNG T. NGUYEN^{2}

1Laboratory of Modelling and Computation, C.N.R.S., University of Grenoble,

B.P. 53x, 38041 Grenoble Cedex, France 2Department of Mathematical Sciences, New Mexico State University, Las Cruces, NM 88003, U.S.A.

(Received October 12, 1991; revised April 20, 1992)

Abstract. This paper concerns the asymptotic validity of the bootstrap method in a non-regular model. Specifically, it is shown that the parametric bootstrap of the change-point parameter in the change-point hazard rate model works.

Key words and phrases: Change-point, consistency of bootstrap method, parametric bootstrapping.